A Complete Bibliography of Publications in the Journal of Software Maintenance: Research and Practice

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

31 January 2018 Version 1.00

Title word cross-reference

2 [CCM94].

0-13-816174-7 [Pay93]. 0-201-63181-4 [Ash95]. 0-412-40520-2 [Pay94b]. 0-412-48440-0 [Dug93]. 0-412-951301 [Inc94b]. 0-442-01361-2 [Gra94a]. 0-471-92940-9 [Pay94a]. 0-471-92985-9 [Yat93]. 0-903969-53-X [Cal91]. 0-971-57011-7 [Inc94a]. 0471-54823-5 [Gra94b].

19.95 [Inc94b]. 1989 [Cha90]. 1991 [NYB92].

23.50 [Gra94b]. 25.00 [Gra94a]. 29.95 [Pay94a].

47.50 [Inc94a].

65 [Pay94b].

ABB [KMFA00]. Abstracting [War93]. Abstraction [KN93, War95a, YLZ00, War95b]. acquisition [Kar92]. actions [Int00]. activities [LH95, MMD93]. Ada [Pay93, BW89]. adapt [Sil94]. Addison [Ash95, Zah99]. adequacy [AS92]. Again [Tra95, Bow93]. aging [Cha94]. Agreement [NLČ99]. Aid [Fer99]. algorithm [HD98, LC95]. Algorithms [MEF99, SW96, KB99b]. Altering [Zim89]. amid [WI97]. among [NLČ99]. Analysis [Tsi96, AZP95, Alk92, CLYH00, DC92, FB98, JLY95, LS98, She97, TG98, TAFM99, TM94, WJBW99]. Andover [Inc94b, Pay94b]. Annealing [MEF99]. Announcement [Ano95a, Ano95b, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, BC94, BC95a, BC95b, BC95c, BC95d, BC95e, BC95f]. Application [GS92, KB99b, TSM96, WOA97, Alk92, FNPQ00, FB98, Leh89, Ric89, SN94]. Applied [MS97, Moh96, Pay94b]. Applying [NLČ99, Tsi96]. Approach [dPLSdP97, CC92, CV00, CLYH00, Gra94b, HH91, Jar94, LH91, LJ94, MOTU93, Ols98, Rin91, Rin92, Sca90, ST00, Sne95, YLZ00, Dug93]. architectural [BFG^{+}00, FATM99]. architecture [AL99, dOBvSdPL98, Can00]. architectures [AOW00]. ART [FATM99]. Articles [WOA97]. aspects [BM90, LS98]. assembly [CHT^{+}90]. Assessing [CCV95, DS98, AS92]. assessment [MB00]. Assessments [MS97]. assistance [Geo92]. Athena [TCM92]. Attitudes [TG98]. attributed [AZP95]. Author [KN95]. Automated [WOA97, vMZ99]. automatic [GS92].

B. [KN95]. back [Bow93]. Base [Cha94, DC92, LG98]. Based [MGF97, AN93, dOBvSdPL98, CHP00, CV00, CLSH00, CK94, EM00, HBK96, HC99, HHKB98, JW98, LH91, MS97, PB94, Sne95, WML91]. basis [Rin92]. batching [KB99b]. be [KAJH99]. behaviour [vMVH97]. benefits [LH90]. better [BM90]. between [Kro95, Sak94]. board [Cha98]. Book [Ash95, Cal91, Cha94, Cha00, Dug93, Gra94b, Gra94a, Inc94a, Inc94b, Pay93, Pay94a, Pay94b, Sea00, Wes96, Yat93, Zah99]. bug [Mar98]. Build [Buf96]. Building [Mar98]. Burlington [Cha94]. Business [HTB^{+}98, MMD93, Kar92, Zah99]. Byme [Pay93].

C [Inc94b, DR96, LC96]. calculation [LC95]. Call [Ano90a]. can [Sil94]. capabilities [BG98]. capability [EM00, Geo92]. Capers [Cha94]. Capturing [Des92]. care [Cha94]. Case [GABG97, TVDF96, Cal90, CLSH00, Hen00, Int00, Mar98, NCS90, Rin91, Rot90]. cased [Cal91]. cases [SB99]. Cem [Gra94a]. Change [GCT^{+}96, AZP95, BW89, Zim89]. Changes [NP90, KB99b, SOSA99]. changing [Sil94]. Chapman [Dug93, Inc94b, Pay94b]. Characteristics [SHH91, Cha93]. Charles [Yat93]. Charts [Haw96]. Chichester [Gra94b, Inc94a, Pay94a, Yat93]. choose [Sak94]. cleanroom [Gra94b]. Client [DL96, Mit95]. Client/Server [DL96, Mit95]. clones [FR99]. cluster [KB99b]. clustering [LG98]. Co [CQ96, Ben92, Cha98]. co-editor [Ben92, Cha98]. Co-operative [CQ96]. COBOL [Cra90, HTB^{+}98, HT99, dPLSdP97, Ves97, LBH93]. COCOMO [HHKB98]. COCOMO-based [HHKB98]. Code [WOA97, Bow93, BL91, CCV95, DC92, FR99, War93, WS95]. Coding [DR96]. Cognitive [Sca90, HH91, HSH99]. Cohesion [NLČ99]. Common [DR96, BM90]. Communication [War95b]. comparative [TG98]. comparison [HC97, PV94]. competitiveness [O’N97]. Compilation [Bow93]. Complex [CHP00, HC99]. complexity [SO95]. component [CLYH00]. comprehension [Sne95]. computer [Gra94a, Leh89]. conceptual [DC92]. Conference [Cha91, Cha90, NYB92]. Configuration [TSM96, CM94, Pay94a]. construction [ST00]. Context [JLY95]. Context-dependent [JLY95]. Control [Haw96, Leh89]. conversion [HT99]. Converting [DL96]. Cooperative [BEM94]. corrective [HC97]. Cost [GABG97, Tsi96, HHKB98]. COTS [CHP00, HC99, HTT97]. COTS-based [CHP00, HC99]. COTS-intensive [HTT97]. Creating [BL91]. Customers [KAJH99]. cycle [Jar94].

D [Pay94a]. Data [NB93, CC92, CL98]. dataflow [JLY95]. David [PC97]. DBT [vMZ99]. Debugging [CL98, HH91, HSH99]. decade [NP90]. Decision [LV95, WML91, Sak94]. Decision-based [WML91]. Decision-driven [LV95]. decisions [Des92]. DECODE [CQ96]. decompilation [Bow93]. decomposition [GS92]. defect [AOW00]. defects [Int00]. defensive [Ves97]. Defining [HBK96]. definition [War95b]. definitions [TAFM99]. Delivery [Moh96, HHKB98]. demand [AN93]. demand-based [AN93]. dependences [LC95]. Dependencies [Buf96]. dependency [AZP95]. dependent [JLY95]. Deriving [AOW00]. Design [Fer99, SI91, Des92, DC92, JW98, Pay93, Rot90, Sil94, Pay94b]. designs [Sca90]. Determinants [KS97]. developing [Rin91]. Development [WOA97, FNPQ00, Kro95, Ric89, WML91, Gra94b]. Dicing [CC97]. difficulties [LH90]. Digital [Pay93]. Discovered [KAJH99]. discussion [KN95, War95a]. Disposable [Voa99]. distinction [Kro95]. Distributed [SW96]. Document [AS92]. documentation [AS92, dOBvSdPL98]. Documentu [dOBvSdPL98]. domain [MMD93]. Dorothy [Ash95]. Downsizing [SN94]. downtime [SF89]. Driven [CDM96, Can00, LV95]. during [vMVH97]. Dyer [Gra94b]. Dynamic [BW89, SF89, WJBW99].

E. [SC95]. early [TM94]. economic [Sch94]. Economics [Sne91]. edition [Gra94a]. editor [Ben92, Cha98, Tra95, War95a].

Editorial [Ano90b, BC89, BC91, Ben92, Cha98]. Education [WB96]. Effective [CL96, Cha93]. effects [SOSA99]. Effort [Lea96, Moh96]. Embedded [TSM96]. Empirical [Fer99, Li97, Alk92, Can00, Jør95, KS97, LH95, SHH91]. engaged [HSH99]. Engineering [Cal91, SW96, Ano90a, BG98, dOBvSdPL98, BL91, Bro93, CC92, CCM94, FATM99, HGG93, Jar94, JW98, KW92, LBH93, LFB95, Leh89, Moh96, MOTU93, Ols98, Sca90, Sne91, YLZ00, YW94]. enhancement [Jon89, Ric89, vMVH97]. Enhancements [Rot90]. Environment [CQ96, FNPQ00, FATM99, Rin91, TCM92, VTP^{+}98]. Estimating [GABG97, Yat93]. European [KCF^{+}91, Ben91]. Evaluating [EM00]. Evaluation [NLČ99, BG98, BKM^{+}98, Can00, HSH99, VTP^{+}98]. Evans [Cal91]. Evolution [Raj97, SB99, KN93, War95a]. Evolutionary [Moh96]. evolving [GS92, MB00]. examination [BCD95, SOSA99]. Example [CL96]. execution [CI91]. exhibited [CL98]. existing [CM94]. experience [OTT^{+}95]. Experimental [NLČ99, SI91]. expert [PK90]. Exploratory [DR96]. exposure [EM00]. Expressions [MGF97]. extraction [HTB^{+}98].

factoring [CC89]. failure [CL98]. Falk [Gra94a]. fault [AOW00, EM00]. Faults [DR96, KAJH99, HC99]. features [WS95]. Fenton [Dug93]. field [SO97]. File [Buf96]. files [AL99]. finder [OYLW94]. finding [LJ94]. fit [DS98]. fitness [DS98]. fitness-for-use [DS98]. fixing [Mar98]. flexible [dOBvSdPL98]. flow [CC92, JLY95]. flow-sensitive [JLY95]. Formal [WB95, CLYH00, LBH93, Rin92]. FORTRAN [HGG93]. foundation [HSH92]. foundations [SB89]. Fourth [KCF^{+}91]. FPA [Yat93]. Framework [VTP^{+}98, AS92, HSH92, KB99a]. frameworks [MB00]. Frank [Cha00, vV00]. free [Cha94]. functional [AD95, Kro95]. Functions [CDM96]. future [Voa99].

generation [AZP95]. Genetic [MEF99]. Geriatric [Cha94]. Gilb [Ash95]. Gillies [Inc94b]. global [LW90, O’N97]. GOTO [MGF97]. government [BCD95]. Graham [Ash95]. graph [CV00]. graph-based [CV00]. graphs [AZP95]. group [Ben91]. grouping [BM90]. guidelines [Zah99]. GUMP [WB96].

Hales [Cal91]. Hall [Dug93, Inc94b, Pay93, Pay94b]. hardback [Pay93, Yat93]. Harmonizing [Ric89]. Hemel [Pay93]. Hempstead [Pay93]. history [AOW00]. Hung [Gra94a]. Hybrid [LC96]. hyperdocuments [HTC95].

iceberg [Tra95]. identification [MOTU93]. Identifying [CDM96, GCT^{+}96]. Impact [GCT^{+}96, FB98, LS98, Sch94, TM94]. Implementing [HBK96]. improvement [Sch93, Zah99]. Improving [Cha00, LFB95, TSM96, Wes96]. Inc. [DL96]. Incremental [HHKB98, Ols98]. Independent [Inc94a]. Index [Leu96, WOA97]. indicators [AS92]. industrial [MB00, OTT^{+}95]. information [Sil94, Voa99]. infrastructure [Mar98]. inquiry [HSH92]. Insights [LW90]. inspection [Ash95]. instrument [SO95]. integrated [Rin91]. Integrating [Sca90]. Integration [Rin92, CLYH00]. Intelligent [Kar92]. intensive [HTT97]. interest [Ben91]. interfaces [YR91]. intermodular [CC92]. interprocedural [JLY95]. Interview [PC97, STR97, SC95]. Inverse [YW94]. Inverse-engineering [YW94]. investigation [KS97, LM94, Rin91, SI91]. ISBN [Ash95, Cal91, Dug93, Gra94b, Gra94a, Inc94a, Inc94b, Pay93, Pay94a, Pay94b, Yat93]. Isolating [HC99]. ISSN [Cha94]. issues [CM94]. Italy [NYB92]. Item [Ben91].

Jack [Gra94a]. Japan [PV94]. John [Gra94b, Inc94a, Pay94a]. Jones [Cha94]. Journal [Ben92]. JSD [BM90]. JSP [Dav90].

Kaner [Gra94a]. KBS [LG98]. Keith [Cal91]. Keller [KN95]. key [YLZ00]. Knowledge [Cha94, LG98, DC92, Kar92, LFB95, MMD93].

languages [VK98]. Large [Moh96, BEM94, FNPQ00, Pay93, SN94, vMVH97]. large-scale [BEM94, vMVH97]. last [NP90]. learning [HH91]. Legacy [CL96, BFG^{+}00, Mit95, WB95]. Letter [Tra95, War95a]. level [BFG^{+}00, Kar90]. Lewis [Inc94a]. libraries [MMD93]. licensing [WI97]. Life [Jar94]. Life-cycle [Jar94]. lines [SB99]. linkage [NB93]. Little [VK98]. Live [Tsi96]. localization [CV00]. logic [Bow93]. London [Cal91, Dug93]. Longman [Zah99]. Ltd [Cal91]. Lurking [Cra90].

M. [War95a]. MA [Cha94]. Macro [Buf96]. MACS [Des92, Geo92]. Mainframe [DL96]. Maintainability [WOA97, Wes96, HTC95, HW91, LH90, SI91]. maintainable [BM90]. Maintaining [BFG^{+}00, Buf96, LC96]. Maintenance [Cha90, Dek92, Fer99, GCT^{+}96, Geo92, GABG97, Haw96, HBK96, HTT97, KCF^{+}91, KDN96, LV95, Lea96, Li97, MS97, Mid95, NYB92, TSM96, TWH96, Tsi96, WB96, AN93, AD95, Alk92, Ben91, Blu95, BKM^{+}98, BCD95, CM94, CHP00, Cha91, Cha93, CHT^{+}90, CI91, Dav90, Des92, DS98, FNPQ00, GS92, HH91, HSH92, HC97, Hen00, HHKB98, Jør95, Kar90, KS97, KTvM^{+}99, Kro95, LH91, LM94, Lea90, LG98, LH95, MMD93, NCS90, NvV00, NP90, O’N97, OYLW94, OTT^{+}95, PB94, PK90, PV94, Ric89, Rin92, Sak94, Sch94, ST00, Sch93, SF89, SHH91, She97, SO95, SO97, SOSA99, SB89, TG98, TML98, Tra95, TM94, VK98, VTP^{+}98, Voa99, YR91, Cha00, Ben92, Sea00]. maintenance-oriented [ST00]. Manage [Haw96, She97]. Management [TSM96, BW89, CM94, Cha93, Mid95, NCS90, NP90, PB94, Sch93, SO97, Inc94b, Pay94a]. map [Mit95]. Mapping [WS95]. maps [NB93]. market [WI97]. Markets [Cal91]. Masthead [Ano89a, Ano89b, Ano90c, Ano90d, Ano90e, Ano90f, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano96k]. Maturity [TSM96]. May [Cha94]. Measure [Can00]. Measure-driven [Can00]. Measurement [AN93, AD95, HBK96, TCM92]. Measurement-based [HBK96]. Measures [NCS90]. Measuring [HTC95, Lea96]. members [Cha98]. Method [GABG97, Moh96, CK94]. Methodologies [TWH96]. Methodology [KDN96, Raj97, LH90]. Methods [Cal91, Lea96, LBH93, WB95, WJBW99, Pay94a]. metric [Sne95]. Metrics [FNPQ00, Fer99, TVDF96, Dug93, HW91, Lea90, LH90, SI91, TCM92]. Michael [Gra94b]. Microcode [TVDF96]. migrating [Mit95]. minimizing [SF89]. MKII [Yat93]. Model [JW98, MS97, CHT^{+}90, HH91, KN93, PB94, Sak94, War95a]. Model-based [JW98, MS97]. Modelling [Kar90, Jon89, Rin91]. models [DS98, WI97]. modified [CK94]. modularization [CCV95]. module [CC89].

Modules [KAJH99]. MOPP [BKM^{+}98]. MSE [Raj97]. multiple [LFB95].

names [AL99]. Nance [KN95]. networks [PB94]. News [Ben91]. Nguyen [Gra94a]. Niessink [Cha00, vV00]. none [Cha94]. Norman [Dug93]. Nostrand [Gra94a]. notion [YLZ00]. numbers [Sne95].

O [Inc94a]. Object

[CV00, GCT^{+}96, Bow93, LH95, LS98, MB00, OYLW94, Pay94b]. Object-orientated [GCT^{+}96]. object-oriented [LH95, LS98, MB00]. objects [LJ94]. observations [SO97]. October [NYB92]. Offshore [KDN96]. old [Sca90]. ontology [KTvM^{+}99]. OO [HLK^{+}97]. operative [CQ96]. Optimal [MEF99]. organization [BCD95]. Organizational [SB89, Sil94]. organizations [BKM^{+}98]. orientated [GCT^{+}96]. oriented [LH95, LS98, MB00, Pay94b, ST00]. OVUM [Cal91].

P [Pay94b]. pages [Cal91]. Paper [Ano90a]. paperback [Ash95, Dug93]. paradox [Blu95]. parallel [HGG93]. Pattern [CLSH00]. Pattern-based [CLSH00]. Peercy [PC97]. Server [DL96, Mit95]. Perception [LM94]. perfective [HC97, Rin92]. PERFORMs [Cra90]. personnel [TG98]. perspective [AN93, NvV00]. Perspectives [Cha00]. Polymorphism [GCT^{+}96]. Porting [dPLSdP97]. pp. [Zah99]. practical [CHP00, Zah99, Sea00]. Practice [Cal90, WOA97, KMFA00]. practices [BCD95, LM94]. precision [TAFM99]. Predicting [HW91]. Prediction [Lea96]. Prentice [Pay93]. Press [Pay93]. pRETS [HGG93]. Preventive [Int00]. Price [Ash95, Cal91, Cha94, Dug93, Gra94b, Inc94a, Inc94b, Pay93, Pay94a, Pay94b, Yat93]. problem [Alk92, KMFA00, WJBW99].

Problems [Dav90]. procedural [CV00]. Procedures [Tsi96]. Process [CDM96, HBK96, WB96, AN93, KMFA00, Mid95, Sch93, Zah99]. Processes [MS97, BKM^{+}98, Can00, HH91, HSH99]. Producers [Leu96]. Producing [BM90]. product [Mid95, SB99]. production [dOBvSdPL98]. Productivity [Cha94]. products [BKM^{+}98]. profiles [KS97]. Progams [DR96]. Program [AZP95, CC97, CQ96, vMVH97, CL98, CHT^{+}90, OYLW94, SF89, Sne95, WS95, YW94]. programmer [HH91]. Programmers [NLČ99, HSH99]. programming [Bow93]. Programs [dPLSdP97, LC96, Bow93, CV00, CK94, GS92, HD98, HTB^{+}98, HT99, Kar92, LJ94, SN94]. Project [GABG97]. Prolog [PK90]. proof [GS92]. proposed [KB99b]. Prosper [Tsi96]. Published [Cal91].

Q [BKM^{+}98]. Q-MOPP [BKM^{+}98]. qualitative [BKM^{+}98]. quality [AS92, Gra94b, HW91, Inc94b, Sch93]. quantitative [HC97, KB99a]. queueing [PB94]. Quoc [Gra94a].

R [Yat93]. R. [KN95]. Rathbone [Cal91]. Re [Moh96, Ano90a, CCM94, Jar94, KW92, Ols98, Sca90, Sne91, CCM94]. Re-engineering [Moh96, Ano90a, CCM94, Jar94, KW92, Ols98, Sca90, Sne91]. reaching [TAFM99]. readability [HTC95]. real [Hen00, YW94]. real-time [Hen00, YW94]. reconnaissance [WS95]. Reconstructive [Li97]. Recovering [AL99]. redevelopment [Sak94]. reduces [HHKB98]. reengineering [CLSH00, YLZ00]. refinement [KN93, War95a]. Regression [Haw96, MEF99, EM00, LW90, vMZ99]. Regular [MGF97]. Reinhold [Gra94a]. releases [SOSA99]. Removal [MGF97]. Removing [FR99]. Reply [KN95]. Report [Cha90, Cha91, KCF^{+}91, NYB92]. reporting [KMFA00]. requirements [Sil94, SOSA99]. Research [Cha94]. residual [Int00]. resolution [KMFA00]. Resolving [Blu95]. restructuring [KB99a]. retrieval [CLYH00]. retroactive [Rot90]. Reusability [PV94]. Reusable [CDM96]. Reuse [Lea96, Li97, AD95, CCM94, Sch94]. revalidating [CK94]. revalidation [HLK^{+}97]. Reverse [BL91, CC92, CCM94, LBH93, SW96, Ano90a, BG98, dOBvSdPL98, Bro93, FATM99, HGG93, JW98, LFB95, MOTU93, YLZ00, Cal91]. Reverse-engineering [BL91, CC92, CCM94, LBH93, dOBvSdPL98, Bro93, HGG93, LFB95, MOTU93]. Review [Ash95, Cal91, Cha94, Cha00, Dug93, Gra94b, Gra94a, Inc94a, Inc94b, Pay93, Pay94a, Pay94b, Sea00, Wes96, Yat93, Zah99, vV00]. rigorous [Dug93]. Risk [Leu96, She97]. road [Mit95]. Robert [Inc94a]. Robinson [Pay94b]. Robotics [KMFA00]. Rock [Cal91]. Rock-Evans [Cal91]. role [CI91]. Rosemary [Cal91]. Routine [Tsi96]. rule [HTB^{+}98]. Rules [NLČ99].

Sano [STR97]. scale [BEM94, FNPQ00, vMVH97]. scavenging [CCV95]. schemes [GS92]. scientific [HSH92]. search [Kar92]. Second [Gra94a]. selection [Hen00]. selective [HLK^{+}97]. semantics [CK94]. semantics-based [CK94]. sensitive [JLY95]. service [NvV00]. Short [Tsi96, War95b]. simple [YW94]. Simulated [MEF99]. simulations [Hen00]. Size [CL96]. sizing [Yat93]. Sleuth [vMZ99]. Slicing [CDM96, HD98]. SOFTM [PK90]. Software [Ben92, Buf96, CM94, Cha90, Cha93, Cha94, Cha00, Dek92, Fer99, GCT^{+}96, GABG97, Haw96, HBK96, Jon89, KMFA00, KCF^{+}91, KAJH99, KDN96, Lea90, Lea96, Leu96, Li97, MS97, NYB92, NvV00, NP90, O’N97, OTT^{+}95, Raj97, Sea00, SO97, TSM96, TWH96, TVDF96, Tsi96, WI97, WS95, WB96, Alk92, AL99, BEM94, Ben91, Blu95, BCD95, Can00, CHP00, Cha91, CLSH00, CI91, DS98, FNPQ00, Geo92, Gra94b, HH91, HSH99, HSH92, HW91, HC97, HLK^{+}97, HTT97, Jar94, Jør95, KB99a, Kar90, KN93, KS97, KTvM^{+}99, KB99b, LM94, Leh89, MMD93, OYLW94, Ols98, PB94, PK90, Pay94a, PV94, Rin92, Sak94, Sch94, ST00, SF89, SHH91, SI91, She97, Sne91, Sne95, SO95, SOSA99, SB99, TML98, TCM92, TM94, Voa99, War95a, WML91, YR91, vMVH97]. software [Cha94, Gra94a, Ash95, Dug93, Inc94b, Pay93, Wes96, Yat93, Zah99]. Sons [Gra94b, Inc94a, Pay94a, Yat93]. Sorrento [NYB92]. Source [WOA97, AL99]. sources [LFB95]. special [Ben91]. Specification [CDM96, LH91, War93]. specification-based [LH91]. Specifications [Bro93, BM90, BL91]. Stability [MB00]. state [KMFA00]. status [Dek92]. steps [CHP00]. strategic [Jar94]. strategies [SO97]. strategy [dOBvSdPL98]. Street [Cal91]. structural [HH91]. structure [MOTU93, OYLW94]. structures [Dav90]. studies [SHH91]. Study [DR96, Fer99, GABG97, Li97, TVDF96, Cal90, CLSH00, HHKB98, Int00, Jør95, LH95, Mar98, NCS90, Rot90]. subsystem [MOTU93]. success [Zah99]. suites [EM00]. support [BFG^{+}00, FB98, MMD93]. survey [SO95]. Swanson [SC95]. symbolic [CI91]. Symons [Yat93]. system [PK90, YLZ00]. Systems [CL96, BEM94, CM94, CHP00, Hen00, HC99, HTT97, Kar90, KW92, LH95, LS98, Mit95, Ols98, Ric89, Sca90, Sil94, Voa99, WB95, Pay93].

T [Cha94]. Takashi [STR97]. task [DS98]. tasks [Jør95]. technique [HLK^{+}97, SF89, TM94]. techniques [BL91, CC89, CCV95, HTB^{+}98, Pay93]. Technology [Pay94b, Can00, DS98]. Tel [Cal91]. teleological [Kar90]. Telogy [DL96]. Test [Hen00, EM00]. Testing [MEF99, LW90, vMZ99, Gra94a]. their [CHP00, EM00]. theoretical [CC92]. Theory [Inc94b, HH91]. thesis [vV00]. time [Hen00, YW94]. Tom [Ash95]. Tool [Buf96, BG98, DS98, VTP^{+}98]. tools [JW98, Pay94a, Cal91]. Toolset [LC96, HGG93]. Total [Sch93]. Traceability [LS98]. Training [TML98]. transaction [KW92]. Transformational [dPLSdP97]. transformations [Rot90]. transitive [LC95]. traps [Ves97]. turbulence [WI97]. Two [CC89, LH95, SB99].

U$1850 [Cal91]. U.S. [BCD95]. UK

[Zah99]. Uncertainty [Leh89].

Understanding [CQ96, Sne95, FB98, OYLW94, SO95, vMVH97]. Unicom [Pay94b]. unstructured [HD98]. updating [SF89]. USA [PV94]. Use [CL96, DS98, LFB95]. user [YR91]. Using [FB98, dPLSdP97, She97, Ves97, AZP95, Bow93, DS98, Rin92, vMZ99]. utilization [DS98].

validation [Inc94a]. Variable [TAFM99]. Variable-precision [TAFM99]. variables [LW90]. verification [Inc94a]. via [LBH93, LH90]. view [AZP95]. volumes [Cal91]. voluminous [CL98]. VSAM [HT99].

W [Pay93]. Ward [War95a]. Wesley [Ash95, Zah99]. Which [KAJH99, Sil94]. Whitgift [Pay94a]. Wiley [Gra94b, Inc94a, Pay94a, Yat93]. will [KAJH99]. Window [YR91]. within [Cra90]. without [Zim89]. work [BEM94, TG98]. Workshop [KCF^{+}91].

X [Cal91].

Y2K [Int00]. year [Mar98, WJBW99].

References

Abran:1995:MFR

[AD95] Alain Abran and Jean-Marc Desharnais. Measurement of functional reuse in maintenance. Journal of Software Maintenance: Research and Practice, 7(4):263–277, July 1995. CODEN JSM-PEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anquetil:1999:RSA

[AL99] Nicolas Anquetil and Timothy C. Lethbridge. Recovering software architecture from the names of source files. Journal of Software Maintenance: Research and Practice, 11(3):201–221, May 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Alkhatib:1992:MPA

[Alk92] Ghazi Alkhatib. The maintenance problem of application software: An empirical analysis. Journal of Software Maintenance: Research and Practice, 4(2):83–104, June 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Abran:1993:MMP

[AN93] Alain Abran and Hong Nguyenkim. Measurement of the maintenance process from a demand-based perspective. Journal of Software Maintenance: Research and Practice, 5(2):63–90, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1989:Ma

[Ano89a]  Anonymous.  Masthead.

Journal of Software Maintenance: Research and Practice, 1(1):fmi, September 1989. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1989:Mb

[Ano89b] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 1(2): fmi, December 1989. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1990:CPR

[Ano90a] Anonymous. Call for paper on reverse engineering and re-engineering. Journal of Software Maintenance: Research and Practice, 2(1): 85, March 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1990:E

[Ano90b] Anonymous. Editorial. Journal of Software Maintenance: Research and Practice, 2(1): 1, March 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1990:Ma

[Ano90c] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 2(1): fmi, March 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1990:Mb

[Ano90d] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 2(2): fmi, June 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1990:Mc

[Ano90e]  Anonymous.  Masthead.

Journal of Software Maintenance: Research and Practice, 2(3):fmi, September 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1990:Md

[Ano90f] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 2(4): fmi, December 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1991:Ma

[Ano91a] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 3(1): fmi, March 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1991:Mb

[Ano91b] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 3(2): fmi, June 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1991:Mc

[Ano91c]  Anonymous.  Masthead.

Journal of Software Maintenance: Research and Practice, 3(3):fmi, September 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1991:Md

[Ano91d] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 3(4): fmi, December 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1992:Ma

[Ano92a] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 4(1): fmi, March 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1992:Mb

[Ano92b] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 4(2): fmi, June 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1992:Mc

[Ano92c]  Anonymous.  Masthead.

Journal of Software Maintenance: Research and Practice, 4(3):fmi, September 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1992:Md

[Ano92d] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 4(4): fmi, December 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1993:Ma

[Ano93a] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 5(1): fmi, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1993:Mb

[Ano93b] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 5(2): fmi, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1993:Mc

[Ano93c] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 5(3): fmi, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1993:Md

[Ano93d] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 5(4): fmi, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1994:Ma

[Ano94a] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 6(1): fmi, January 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1994:Mb

[Ano94b] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 6(2): fmi, March 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1994:Mc

[Ano94c] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 6(3): fmi, May 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1994:Md

[Ano94d] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 6(4): fmi, July 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1994:Me

[Ano94e]  Anonymous.  Masthead.

Journal of Software Maintenance: Research and Practice, 6(5):fmi, September 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1994:Mf

[Ano94f] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 6(6): fmi, November 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1995:Aa

[Ano95a] Anonymous. Announcement.

Journal of Software Maintenance: Research and Practice, 7(1):i, January 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1995:Ab

[Ano95b] Anonymous. Announcement.

Journal of Software Maintenance: Research and Practice, 7(2):i, March 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1995:Ma

[Ano95c] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 7(1): fmi, January 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1995:Mb

[Ano95d] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 7(2): fmi, March 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1995:Mc

[Ano95e] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 7(3): fmi, May 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1995:Md

[Ano95f] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 7(4): fmi, July 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1995:Me

[Ano95g]  Anonymous.  Masthead.

Journal of Software Maintenance: Research and Practice, 7(5):fmi, September 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1995:Mf

[Ano95h] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 7(6): fmi, November 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1996:Aa

[Ano96a] Anonymous. Announcement.

Journal of Software Maintenance: Research and Practice, 8(2):i, March 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1996:Ab

[Ano96b]  Anonymous.  Announce-

ment. Journal of Software Maintenance: Research and Practice, 8(3):i, May 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1996:Ac

[Ano96c]  Anonymous.  Announce-

ment. Journal of Software Maintenance: Research and Practice, 8(4):i, July 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1996:Ad

[Ano96d] Anonymous. Announcement. Journal of Software Maintenance: Research and Practice, 8(5):fmi, September 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1996:Ae

[Ano96e] Anonymous. Announcement. Journal of Software Maintenance: Research and Practice, 8(6):fmi, November 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1996:Ma

[Ano96f] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 8(1): fmi, January 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1996:Mb

[Ano96g] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 8(2): fmi, March 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1996:Mc

[Ano96h] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 8(3): fmi, May 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1996:Md

[Ano96i] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 8(4): fmi, July 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1996:Me

[Ano96j]  Anonymous.  Masthead.

Journal of Software Maintenance: Research and Practice, 8(5):fmi, September 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Anonymous:1996:Mf

[Ano96k] Anonymous. Masthead. Journal of Software Maintenance: Research and Practice, 8(6): fmi, November 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Andrews:2000:DFA

[AOW00] Anneliese Amschler Andrews, Magnus C. Ohlsson, and Claes Wohlin. Deriving fault architectures from defect history. Journal of Software Maintenance: Research and Practice, 12(5): 287–304, September 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Arthur:1992:DQI

[AS92] James D. Arthur and K. Todd Stevens. Document quality indicators: A framework for assessing documentation adequacy. Journal of Software Maintenance: Research and Practice, 4(3): 129–142, September 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Ashley:1995:BRS

[Ash95] John Ashley. Book review: Software inspection, Tom Gilb and Dorothy Graham, Addison Wesley, 1993 ISBN 0-201-63181-4. Price £24.95 (paperback). Journal of Software Maintenance: Research and Practice, 7(3):221– 222, May 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Al-Zoubi:1995:PVG

[AZP95] Ratib Al-Zoubi and Atul Prakash. Program view generation and change analysis using attributed dependency graphs. Journal of Software Maintenance: Research and Practice, 7(4):239– 261, July 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bennet:1989:E

[BC89] Keith Bennet and Mel Colter.

Editorial. Journal of Software Maintenance: Research and Practice, 1(1):1, September 1989. CODEN JSM-PEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bennet:1991:E

[BC91] Keith Bennet and Mel Colter.

Editorial. Journal of Software Maintenance: Research and Practice, 3(2):63, June 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bennett:1994:A

[BC94] Keith Bennett and Ned Chapin. Announcement. Journal of Software Maintenance: Research and Practice, 6(6):285, November 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bennett:1995:Aa

[BC95a] Keith Bennett and Ned Chapin. Announcement. Journal of Software Maintenance: Research and Practice, 7(1):1, January 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bennett:1995:Ab

[BC95b] Keith Bennett and Ned Chapin. Announcement.

Journal of Software Maintenance: Research and Practice, 7(2):75, March 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bennett:1995:Ac

[BC95c] Keith Bennett and Ned Chapin. Announcement. Journal of Software Maintenance: Research and Practice, 7(3):149, May 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bennett:1995:Ad

[BC95d] Keith Bennett and Ned Chapin. Announcement. Journal of Software Maintenance: Research and Practice, 7(4):i, July 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bennett:1995:Ae

[BC95e] Keith Bennett and Ned Chapin. Announcement. Journal of Software Maintenance: Research and Practice, 7(5):301, September 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bennett:1995:Af

[BC95f] Keith Bennett and Ned Chapin. Announcement. Journal of Software Maintenance: Research and Practice, 7(6):ii, November 1995.

CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Brown:1995:ESM

[BCD95] Alan W. Brown, Alan M.

Christie, and Susan A. Dart. An examination of software maintenance practices in a U.S. government organization. Journal of Software Maintenance: Research and Practice, 7(4):223–238, July 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Belkhatir:1994:CWL

[BEM94] Noureddine Belkhatir, Jacky Estublier, and Walcélio L. Melo. Cooperative work in large-scale software systems. Journal of Software Maintenance: Research and Practice, 6(6):319–335, November 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bennett:1991:NIE

[Ben91] Keith H. Bennett. News item: European special interest group in software maintenance. Journal of Software Maintenance: Research and Practice, 3(1):59–61, March 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bennett:1992:ENC

[Ben92] Keith Bennett. Editorial: New co-editor for Journal of Software Maintenance. Journal of Software Maintenance: Research and Practice, 4(4):181, December 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bril:2000:MLT

[BFG^{+}00] Reinder J. Bril, Loe M. G. Feijs, André Glas, René L. Krikhaar, and M. (Thijs) R. M. Winter. Maintaining a legacy: towards support at the architectural level. Journal of Software Maintenance: Research and Practice, 12(3):143–170, May 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bellay:1998:ERE

[BG98] Berndt Bellay and Harald Gall. An evaluation of reverse engineering tool capabilities. Journal of Software Maintenance: Research and Practice, 10(5):305–331, September 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Briand:1998:QMQ

[BKM^{+}98] Lionel Briand, Yong-Mi Kim, Walcélio Melo, Carolyn Seaman, and Victor R. Basili. Q-MOPP: qualitative evaluation of maintenance organizations, processes and products. Journal of Software Maintenance: Research and Practice, 10(4):249–278, July 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Breuer:1991:CSC

[BL91] P. T. Breuer and K. Lano.

Creating specifications from code: Reverse-engineering techniques. Journal of Software Maintenance: Research and Practice, 3(3):145–162, September 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Blum:1995:RSM

[Blu95]  Bruce I. Blum.  Resolv-

ing the software maintenance paradox. Journal of Software Maintenance: Research and Practice, 7(1):3– 26, January 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Borgers:1990:PBM

[BM90] M. Borgers and M. Munro.

Producing better maintainable JSD specifications by grouping common aspects. Journal of Software Maintenance: Research and Practice, 2(1):61–80, March 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Bowen:1993:POC

[Bow93] Jonathan Bowen. From programs to object code and back again using logic programming: Compilation and decompilation. Journal of Software Maintenance: Research and Practice, 5(4):205– 234, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Brown:1993:SRE

[Bro93] Alan J. Brown. Specifications and reverse-engineering. Journal of Software Maintenance: Research and Practice, 5(3):147–153, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Buffenbarger:1996:STM

[Buf96] Jim Buffenbarger. A software tool for maintaining file and macro build dependencies. Journal of Software Maintenance: Research and Practice, 8(6):421–431, November 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Burns:1989:DCM

[BW89] A. Burns and A. J. Wellings. Dynamic change management and Ada. Journal of Software Maintenance: Research and Practice, 1(2): 121–131, December 1989.

CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Calow:1990:PCS

[Cal90] H. Calow. Practice case study.

Journal of Software Maintenance: Research and Practice, 2(1):81–84, March 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Calow:1991:BRR

[Cal91]  Hilary Calow.  Book re-

view: Reverse Engineering: Markets, Methods and Tools, Rosemary Rock-Evans and Keith Hales. Published by OVUM Ltd (7 Rathbone Street, London. Tel: 071 255 2670), 1990. ISBN 0-903969-53-X, 750 pages, 2 volumes. Price: £995 (U$1850) cased. Journal of Software Maintenance: Research and Practice, 3(1):57–58, March 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Cantone:2000:MDP

[Can00] Giovanni Cantone. Measuredriven processes and architecture for the empirical evaluation of software technology. Journal of Software Maintenance: Research and Practice, 12(1):47–78, January 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Calliss:1989:TMF

[CC89] F. W. Calliss and B. J. Cornelius. Two module factoring techniques. Journal of Software Maintenance: Research and Practice, 1 (2):81–89, December 1989. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Canfora:1992:REI

[CC92] G. Canfora and A. Cimitile. Reverse-engineering and intermodular data flow: A theoretical approach. Journal of Software Maintenance: Research and Practice, 4(1):37– 59, March 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chen:1997:PD

[CC97] T. Y. Chen and Y. Y. Cheung. On program dicing. Journal of Software Maintenance: Research and Practice, 9(1):33–46, January 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Canfora:1994:RRE

[CCM94] G. Canfora, A. Cimitile, and M. Munro. RE^{2}: Reverseengineering and reuse reengineering. Journal of Software Maintenance: Research and Practice, 6(2):53– 72, March 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Canfora:1995:AMC

[CCV95] G. Canfora, A. Cimitile, and G. Visaggio. Assessing modularization and code scavenging techniques. Journal of Software Maintenance: Research and Practice, 7(5): 317–331, September 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Cimitile:1996:SDS

[CDM96] Aniello Cimitile, Andrea De Lucia, and Malcolm Munro. A specification driven slicing process for identifying reusable functions. Journal of Software Maintenance: Research and Practice, 8(3):145– 178, May 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chapin:1990:RCS

[Cha90] Ned Chapin. Report: Conference on Software Maintenance — 1989: a report. Journal of Software Maintenance: Research and Practice, 2(2):143–144, June 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chapin:1991:RCS

[Cha91] Ned Chapin. Report: Conference on software maintenance — 1990: A report.

Journal of Software Maintenance: Research and Practice, 3(2):125–127, June 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chapin:1993:SMC

[Cha93] Ned Chapin. Software maintenance characteristics and effective management. Journal of Software Maintenance: Research and Practice, 5(2): 91–100, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chapin:1994:BRG

[Cha94] Dr. Ned Chapin. Book review: Geriatric care for aging software, T. Capers Jones in Knowledge Base 1, 1 (May 1992). Software Productivity Research Inc, Burlington, MA. ISSN none. Price: free. Journal of Software Maintenance: Research and Practice, 6(3):161–162, May 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chapin:1998:ENC

[Cha98] Ned Chapin. Editorial: New co-editor and board members. Journal of Software Maintenance: Research and Practice, 10(1):1, January 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chapin:2000:BRR

[Cha00] Ned Chapin. Book review:

Review of Perspectives on Improving Software Maintenance by Frank Niessink. Journal of Software Maintenance: Research and Practice, 12(3):187–195, May 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Carney:2000:CCB

[CHP00] David Carney, Scott A.

Hissam, and Daniel Plakosh. Complex COTS-based software systems: practical steps for their maintenance. Journal of Software Maintenance: Research and Practice, 12 (6):357–376, November 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chen:1990:MAP

[CHT^{+}90] S. Chen, K. G. Heisler, W. T.

Tsai, X. Chen, and E. Leung. A model for assembly program maintenance. Journal of Software Maintenance: Research and Practice, 2(1):3–32, March 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Coward:1991:RSE

[CI91] P. D. Coward and D. C. Ince.

The role of symbolic execution in software maintenance. Journal of Software Maintenance: Research and Practice, 3(4):183–192, December 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chung:1994:SBM

[CK94] In Sang Chung and Yong Rae Kwon. A semantics-based method for revalidating modified programs. Journal of Software Maintenance: Research and Practice, 6(1):15– 33, January 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chapin:1996:ESE

[CL96] Ned Chapin and Tony S. Lau. Effective size: An example of use from legacy systems. Journal of Software Maintenance: Research and Practice, 8(2):101–116, March 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chan:1998:DPF

[CL98] Tat W. Chan and Arun Lakhotia. Debugging program failure exhibited by voluminous data. Journal of Software Maintenance: Research and Practice, 10 (2):111–150, March 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chu:2000:PBS

[CLSH00] William C. Chu, Chih-Wei Lu, Chih-Peng Shiu, and Xudong He. Pattern-based software reengineering: a case study. Journal of Software Maintenance: Research and Practice, 12(2):121–141, March 2000. CODEN JSM-PEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chu:2000:FAC

[CLYH00] William C. Chu, Chih-Wei Lu, Hongji Yang, and Xudong He. A formal approach for component retrieval and integration analysis. Journal of Software Maintenance: Research and Practice, 12 (6):325–342, November 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Capretz:1994:SCM

[CM94] Miriam A. M. Capretz and Malcolm Munro. Software configuration management issues in the maintenance of existing systems. Journal of Software Maintenance: Research and Practice, 6(1):1– 14, January 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Chin:1996:DCO

[CQ96] David N. Chin and Alex Quilici. DECODE: A cooperative program understanding environment. Journal of Software Maintenance: Research and Practice, 8(1):

3–33, January 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Crawford:1990:LWC

[Cra90] M. A. Crawford. Lurking within COBOL PERFORMs. Journal of Software Maintenance: Research and Practice, 2(1):33–46, March 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Carver:2000:OLP

[CV00] Doris L. Carver and Ramachenga Valasareddi. Object localization in procedural programs: a graphbased approach. Journal of Software Maintenance: Research and Practice, 12(5): 305–323, September 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Davies:1990:PMJ

[Dav90] C. G. Davies. Problems of maintenance of JSP structures. Journal of Software Maintenance: Research and Practice, 2(3):175–192, September 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Dietrich:1992:CDC

[DC92] Suzanne W. Dietrich and Frank W. Calliss. A conceptual design for a code analysis knowledge base. Journal of Software Maintenance: Research and Practice, 4(1): 19–36, March 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Dekleva:1992:SMS

[Dek92] Sasa M. Dekleva. Software maintenance: 1990 status. Journal of Software Maintenance: Research and Practice, 4(4):233–247, December 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Desclaux:1992:CDM

[Des92] Christine Desclaux. Capturing design and maintenance decisions with MACS. Journal of Software Maintenance: Research and Practice, 4(4):215–231, December 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Duncan:1996:CMC

[DL96] Doris G. Duncan and Sateesh B.

Lele. Converting from mainframe to client/server at Telogy Inc. Journal of Software Maintenance: Research and Practice, 8(5):321–344, September 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Braga:1998:DFA

[dOBvSdPL98] Christiano de Oliveira Braga, Arndt von Staa, and Julio Cesar Sampaio do Prado Leite. Documentu: a flexible architecture for documentation production based on a reverse-engineering strategy. Journal of Software Maintenance: Research and Practice, 10(4):279–303, July 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Leite:1997:PCP

[dPLSdP97] Julio Cesar Sampaio do Prado Leite, Marcelo Sant’Anna, and Antonio Francisco do Prado. Porting COBOL programs using a transformational approach. Journal of Software Maintenance: Research and Practice, 9(1):3–31, January 1997. CODEN JSM-PEU. ISSN 1040-550X (print), 1096-908X (electronic).

Duncan:1996:ESC

[DR96] Ishbel Duncan and Dave Robson. An exploratory study of common coding faults in C progams. Journal of Software Maintenance: Research and Practice, 8(4):241– 256, July 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Dishaw:1998:ASM

[DS98] Mark T. Dishaw and Diane M. Strong. Assessing software maintenance tool utilization using task–technology fit and fitness-for-use models. Journal of Software Maintenance: Research and Practice, 10(3):151–179, May 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Dugard:1993:BRS

[Dug93] Pat Dugard. Book review:

Software metrics: A rigorous approach, Norman E. Fenton. Chapman & Hall (London) 1991. ISBN 0-412-48440-0. Price £19.95 (paperback). Journal of Software Maintenance: Research and Practice, 5(1):59, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Elbaum:2000:ERT

[EM00] Sebastian G. Elbaum and John C. Munson. Evaluating regression test suites based on their fault exposure capability. Journal of Software Maintenance: Research and Practice, 12(3): 171–184, May 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Fiutem:1999:AAR

[FATM99] Roberto Fiutem, Giulio Antoniol, Paolo Tonella, and Ettore Merlo. ART: an architectural reverse engineering environment. Journal of Software Maintenance: Research and Practice, 11(5): 339–364, September 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Fyson:1998:UAU

[FB98] M. Joanna Fyson and Cornelia Boldyreff. Using application understanding to support impact analysis. Journal of Software Maintenance: Research and Practice, 10(2): 93–110, March 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Ferneley:1999:DMA

[Fer99] Elaine H. Ferneley. Design metrics as an aid to software maintenance: An empirical study. Journal of Software Maintenance: Research and Practice, 11(1):55– 72, January 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Fasolino:2000:MDM

[FNPQ00] Anna Rita Fasolino, Domenico Natale, Alessio Poli, and Alessandro Alberigi Quaranta. Metrics in the development and maintenance of software: an application in a large scale environment. Journal of Software Maintenance:

Research and Practice, 12 (6):343–355, November 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Fanta:1999:RCC

[FR99] Richard Fanta and Václav Rajlich. Removing clones from the code. Journal of Software Maintenance: Research and Practice, 11(4): 223–243, July 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Granja-Alvarez:1997:MEM

[GABG97] Juan Carlos Granja-Alvarez and Manuel José Barranco-García. A method for estimating maintenance cost in a software project: A case study. Journal of Software Maintenance: Research and Practice, 9(3):161–175, May 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Gao:1996:IPC

[GCT^{+}96] Jerry Gao, Chris Chen, Y. Toyoshima, David Kung, and Pei Hsia. Identifying polymorphism change and impact in object-orientated software maintenance. Journal of Software Maintenance: Research and Practice, 8 (6):357–387, November 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Georges:1992:MMA

[Geo92] Mari Georges. MACS: Maintenance assistance capability for software. Journal of Software Maintenance: Research and Practice, 4(4): 199–213, December 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Graham:1994:BRT

[Gra94a] Dorothy R. Graham. Book review: Testing computer software, Cem Kaner, Jack Falk and Hung Quoc Nguyen, Second edition, Van Nostrand Reinhold, 1993. ISBN 0-442-01361-2. £25.00. Journal of Software Maintenance: Research and Practice, 6(3):162– 163, May 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Graham:1994:BRC

[Gra94b] Dorothy R. Graham. Book review: The cleanroom approach to software quality development, Michael Dyer. John Wiley & Sons, Chichester, 1992, ISBN 0471-54823-5. Price: £23.50. Journal of Software Maintenance: Research and Practice, 6 (2):100–102, March 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Gopal:1992:AAD

[GS92] R. Gopal and S. R. Schach.

Application of automatic decomposition schemes in proof maintenance for evolving programs. Journal of Software Maintenance: Research and Practice, 4(4):183–198, December 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Haworth:1996:RCC

[Haw96] Dwight A. Haworth. Regression control charts to manage software maintenance. Journal of Software Maintenance: Research and Practice, 8(1):35–48, January 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Henry:1996:DIM

[HBK96] Joel Henry, Robert Blasewitz, and David Kettinger. Defining and implementing a measurement-based software maintenance process. Journal of Software Maintenance: Research and Practice, 8(2):79–100, March 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Henry:1997:QCP

[HC97] Joel E. Henry and James P.

Cain. A quantitative comparison of perfective and corrective software maintenance. Journal of Software Maintenance: Research and Practice, 9(5):281–297, September 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Hissam:1999:IFC

[HC99] Scott A. Hissam and David Carney. Isolating faults in complex COTS-based systems. Journal of Software Maintenance: Research and Practice, 11(3):183–199, May 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Harman:1998:NAS

[HD98] Mark Harman and Sebastian Danicic. A new algorithm for slicing unstructured programs. Journal of Software Maintenance: Research and Practice, 10(6): 415–441, November 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Henry:2000:TCS

[Hen00] Joel E. Henry. Test case selection for simulations in the maintenance of realtime systems. Journal of Software Maintenance: Research and Practice, 12(4): 229–248, July 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Harrison:1993:PPR

[HGG93] Warren Harrison, Cynthia Gens, and Bruce Gifford. pRETS: A parallel reverseengineering toolset for FOR-TRAN. Journal of Software Maintenance: Research and Practice, 5(1):37–57, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Hale:1991:TMP

[HH91] David P. Hale and Dwight A.

Haworth. Towards a model of programmers’ cognitive processes in software maintenance: A structural learning theory approach for debugging. Journal of Software Maintenance: Research and Practice, 3(2):85–106, June 1991. CODEN JSM-PEU. ISSN 1040-550X (print), 1096-908X (electronic).

Hsia:1998:IDR

[HHKB98] Pei Hsia, Chih-Tung Hsu, David Chenho Kung, and Eric J. Byrne. Incremental delivery reduces maintenance cost: a COCOMObased study. Journal of Software Maintenance: Research and Practice, 10(4): 225–247, July 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Hsia:1997:TSR

[HLK^{+}97] Pei Hsia, Xiaolin Li, David Chenho Kung, Chih-Tung Hsu, Liang Li, Yasufumi Toyoshima, and Cris Chen. A technique for the selective revalidation of OO software. Journal of Software Maintenance: Research and Practice, 9(4):217– 233, July 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Haworth:1992:FSM

[HSH92] Dwight A. Haworth, Shane Sharpe, and David P. Hale. A framework for software maintenance: A foundation for scientific inquiry. Journal of Software Maintenance: Research and Practice, 4(2):105–117, June 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Hale:1999:ECP

[HSH99] Joanne E. Hale, Shane Sharpe, and David P. Hale. An evaluation of the cognitive processes of programmers engaged in software debugging. Journal of Software Maintenance: Research and Practice, 11(2):73–91, March 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Huang:1999:VCC

[HT99] Hai Huang and Wei-Tek Tsai. VSAM conversion for COBOL programs. Journal of Software Maintenance: Research and Practice, 11 (5):311–337, September 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Huang:1998:BRE

[HTB^{+}98] Hai Huang, Wei-Tek Tsai, Sourav Bhattacharya, Xiaoping Chen, Yamin Wang, and Jianhua Sun. Business rule extraction techniques for COBOL programs. Journal of Software Maintenance: Research and Practice, 10(1):3–35, January 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Hatzimanikatis:1995:MRM

[HTC95] A. E. Hatzimanikatis, C. T.

Tsalidis, and D. Christodoulakis. Measuring the readability and maintainability of hyperdocuments. Journal of Software Maintenance: Research and Practice, 7(2):77– 90, March 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Hybertson:1997:MCI

[HTT97] Duane W. Hybertson, Anh D.

Ta, and William M. Thomas. Maintenance of COTS-intensive software systems. Journal of Software Maintenance: Research and Practice, 9(4):203– 216, July 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Henry:1991:PMS

[HW91] Sallie Henry and Steve Wake. Predicting maintainability with software quality metrics. Journal of Software Maintenance: Research and Practice, 3(3):129–143, September 1991. CODEN JSM-PEU. ISSN 1040-550X (print), 1096-908X (electronic).

Ince:1994:BRI

[Inc94a] Professor Darrel Ince. Book review: Independent verification and validation, Robert O. Lewis. John Wiley & Sons, Chichester, 1992, ISBN 0-971-57011-7. Price: £47.50. Journal of Software Maintenance: Research and Practice, 6(2):99, March 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Ince:1994:BRS

[Inc94b] Professor Darrel Ince. Book review: Software quality: Theory and management, A. C. Gillies. Chapman & Hall, Andover, 1992, ISBN 0-412-951301. Price: £19.95. Journal of Software Maintenance: Research and Practice, 6(2):99–100, March 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Interesse:2000:PAR

[Int00] Michelangelo Interesse. Preventive actions for residual Y2K defects: a case study. Journal of Software Maintenance: Research and Practice, 12(1):1–23, January 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Jarzabek:1994:LCA

[Jar94] Stan Jarzabek. Life-cycle approach to strategic reengineering of software. Journal of Software Maintenance: Research and Practice, 6 (6):287–317, November 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Johmann:1995:CDF

[JLY95] Kurt R. Johmann, Sying-Syang Liu, and Stephen S. Yau. Context-dependent flow-sensitive interprocedural dataflow analysis. Journal of Software Maintenance: Research and Practice, 7(3):177– 202, May 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Jones:1989:SEM

[Jon89]  C. Jones.  Software en-

hancement modelling. Journal of Software Maintenance: Research and Practice, 1 (2):91–100, December 1989. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Jorgensen:1995:ESS

[Jør95] Magne Jørgensen. An empirical study of software maintenance tasks. Journal of Software Maintenance: Research and Practice, 7(1):27– 48, January 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Jarzabek:1998:MBD

[JW98] Stan Jarzabek and Guosheng Wang. Model-based design of reverse engineering tools. Journal of Software Maintenance: Research and Practice, 10(5):353–380, September 1998. CODEN JSM-PEU. ISSN 1040-550X (print), 1096-908X (electronic).

Khoshgoftaar:1999:WSM

[KAJH99] Taghi M. Khoshgoftaar, Edward B. Allen, Wendell D. Jones, and John P. Hudepohl. Which software modules have faults which will be discovered by customers? Journal of Software Maintenance: Research and Practice, 11(1):1–18, January 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Karakostas:1990:MMS

[Kar90] V. Karakostas. Modelling and maintenance software systems at the teleological level. Journal of Software Maintenance: Research and Practice, 2(1):47–59, March 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Karakostas:1992:ISA

[Kar92] V. Karakostas. Intelligent search and acquisition of business knowledge from programs. Journal of Software Maintenance: Research and Practice, 4(1):1–17, March 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Kang:1999:QFS

[KB99a] Byung-Kyoo Kang and James M.

Bieman. A quantitative framework for software restructuring. Journal of Software Maintenance: Research and Practice, 11(4): 245–284, July 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Krohn:1999:ACA

[KB99b] Uwe Krohn and Cornelia Boldyreff. Application of cluster algorithms for batching of proposed software changes. Journal of Software Maintenance: Research and Practice, 11(3):151–165, May 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Kenning:1991:RRF

[KCF^{+}91] Rachel J. Kenning, Simon D. Cooper, Robert M. Freeman, David S. Hinley, and Malcolm Munro. Report: Report on the Fourth European Software Maintenance Workshop. Journal of Software Maintenance: Research and Practice, 3(1):45–56, March 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Kumar:1996:OSM

[KDN96] M. Pavan Kumar, V. Sita Rama Das, and N. Netaji. Offshore software maintenance methodology. Journal of Software Maintenance: Research and Practice, 8(3):179– 197, May 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Kajko-Mattsson:2000:SPR

[KMFA00] Mira Kajko-Mattsson, Stefan Forssander, and Gunnar Andersson. Software problem reporting and resolution process at ABB Robotics AB: state of practice. Journal of Software Maintenance: Research and Practice, 12 (5):255–285, September 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Keller:1993:ARM

[KN93] Benjamin J. Keller and Richard E. Nance. Abstraction refinement: A model of software evolution. Journal of Software Maintenance: Research and Practice, 5 (3):123–145, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic). See discussion [War95a] and reply [KN95].

Keller:1995:ARR

[KN95] Ben Keller and Richard Nance. Authors’ reply: Reply to discussion by B. Keller and R. Nance. Journal of Software Maintenance: Research and Practice, 7(5): 379–380, September 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic). See [KN93].

Krogstie:1995:DBF

[Kro95] John Krogstie. On the distinction between functional development and functional maintenance. Journal of Software Maintenance: Research and Practice, 7(6): 383–403, November 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Kemerer:1997:DSM

[KS97] Chris F. Kemerer and Sandra A. Slaughter. Determinants of software maintenance profiles: an empirical investigation. Journal of Software Maintenance: Research and Practice, 9(4):235– 251, July 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Kitchenham:1999:TOS

[KTvM^{+}99] Barbara A. Kitchenham, Guilherme H. Travassos, Anneliese von Mayrhauser, Frank Niessink, Norman F. Schneidewind, Janice Singer, Shingo Takada, Risto Vehvilainen, and Hongji Yang. Towards an ontology of software maintenance. Journal of Software Maintenance: Research and Practice, 11 (6):365–389, November 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Kozaczynski:1992:RET

[KW92] Wojtek Kozaczynski and Norman Wilde. On the reengineering of transaction systems. Journal of Software Maintenance: Research and Practice, 4(3):143–162, September 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Lano:1993:REC

[LBH93] K. Lano, P. T. Breuer, and H. Haughton. Reverseengineering Cobol via formal methods. Journal of Software Maintenance: Research and Practice, 5(1):13– 35, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Livadas:1995:NAC

[LC95] Panos E. Livadas and Stephen Croll. A new algorithm for the calculation of transitive dependences. Journal of Software Maintenance: Research and Practice, 7(3):151– 176, May 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Linos:1996:TMH

[LC96] Panagiotis K. Linos and Vincent Courtois. A toolset for maintaining hybrid C++ programs. Journal of Software Maintenance: Research and Practice, 8(6):389–419, November 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Leach:1990:SMS

[Lea90]  Ronald J. Leach.  Soft-

ware metrics and software maintenance. Journal of Software Maintenance: Research and Practice, 2(2):133– 142, June 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Leach:1996:MMS

[Lea96] Ronald J. Leach. Methods of measuring software reuse for the prediction of maintenance effort. Journal of Software Maintenance: Research and Practice, 8(5): 309–320, September 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Lehman:1989:UCA

[Leh89]  M. M. Lehman.  Uncer-

tainty in computer application and its control through the engineering of software. Journal of Software Maintenance: Research and Practice, 1(1):3–27, September 1989. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Leung:1996:RIS

[Leu96] Hareton K. N. Leung. A risk index for software producers. Journal of Software Maintenance: Research and Practice, 8(5):281–294, September 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Layzell:1995:IRE

[LFB95] P. J. Layzell, M. J. Freeman, and P. Benedusi. Improving reverse-engineering through the use of multiple knowledge sources. Journal of Software Maintenance: Research and Practice, 7(4):279– 299, July 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Lee:1998:KBC

[LG98] Ook Lee and Paul Gray.

Knowledge base clustering for KBS maintenance. Journal of Software Maintenance: Research and Practice, 10 (6):395–414, November 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Lewis:1990:BDM

[LH90] John A. Lewis and Sallie M.

Henry. On the benefits and difficulties of a maintainability via metrics methodology. Journal of Software Maintenance: Research and Practice, 2(2):113–131, June 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Lano:1991:SBA

[LH91] K. Lano and H. Haughton. A specification-based approach to maintenance. Journal of Software Maintenance: Research and Practice, 3(4): 193–213, December 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Li:1995:ESM

[LH95] Wei Li and Sallie Henry. An empirical study of maintenance activities in two objectoriented systems. Journal of Software Maintenance: Research and Practice, 7(2):131– 147, March 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Li:1997:ESS

[Li97] Wei Li. An empirical study of software reuse in reconstructive maintenance. Journal of Software Maintenance: Research and Practice, 9(2): 69–83, March 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Livadas:1994:NAF

[LJ94] Panos E. Livadas and Theodore Johnson. A new approach to finding objects in programs. Journal of Software Maintenance: Research and Practice, 6(5):249–260, September 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Layzell:1994:ISM

[LM94] Paul J. Layzell and Linda A. Macaulay. An investigation into software maintenance — perception and practices. Journal of Software Maintenance: Research and Practice, 6(3):105–120, May 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Lindvall:1998:TAI

[LS98] Mikael Lindvall and Kristian Sandahl. Traceability aspects of impact analysis in object-oriented systems. Journal of Software Maintenance: Research and Practice, 10(1):37–57, January 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Lanubile:1995:DDM

[LV95] Filippo Lanubile and Giuseppe Visaggio. Decision-driven maintenance. Journal of Software Maintenance: Research and Practice, 7(2):91– 115, March 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Leung:1990:ITR

[LW90] Hareton K. N. Leung and Lee White. Insights into testing and regression testing global variables. Journal of Software Maintenance: Research and Practice, 2 (4):209–222, December 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Marcoccia:1998:BIF

[Mar98] Louis J. Marcoccia. Building infrastructure for fixing the year 2000 bug: a case study. Journal of Software Maintenance: Research and Practice, 10(5): 333–352, September 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Mattsson:2000:SAE

[MB00] Michael Mattsson and Jan Bosch. Stability assessment of evolving industrial object-oriented frameworks. Journal of Software Maintenance: Research and Practice, 12(2):79–102, March 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Mansour:1999:SAG

[MEF99] Nashat Mansour and Khalid El-Fakih. Simulated annealing and genetic algorithms for optimal regression testing. Journal of Software Maintenance: Research and Practice, 11(1):19–34, January 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Morris:1997:GRB

[MGF97] Paul H. Morris, Ronald A. Gray, and Robert E. Filman. GOTO removal based on regular expressions. Journal of Software Maintenance: Research and Practice, 9(1):47– 66, January 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Middleton:1995:MMP

[Mid95] Peter Middleton. Maintenance management: From product to process. Journal of Software Maintenance: Research and Practice, 7(1):63– 73, January 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Mittra:1995:RMM

[Mit95] Sitansu S. Mittra. A road map for migrating legacy systems to client/server. Journal of Software Maintenance: Research and Practice, 7 (2):117–130, March 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Mendes-Moreira:1993:BDK

[MMD93] H. M. C. L. Mendes-Moreira and C. G. Davies. Business domain knowledge libraries to support software maintenance activities. Journal of Software Maintenance: Research and Practice, 5 (3):165–179, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Mohan:1996:EMD

[Moh96]  Shantha Mohan.  Evolu-

tionary method of delivery as applied to a large reengineering effort. Journal of Software Maintenance: Research and Practice, 8(4):269– 280, July 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Muller:1993:REA

[MOTU93] Hausi A. Müller, Mehmet A.

Orgun, Scott R. Tilley, and James S. Uhl. A reverseengineering approach to subsystem structure identification. Journal of Software Maintenance: Research and Practice, 5(4):181–204, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Masengale:1997:MBA

[MS97] Derick E. Masengale and Michael J. Schrank. Modelbased assessments applied to software maintenance processes. Journal of Software Maintenance: Research and Practice, 9(2):85–101, March 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Neil:1993:DLM

[NB93] Martin Neil and Richard Bache. Data linkage maps. Journal of Software Maintenance: Research and Practice, 5(3):155–164, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Neil:1990:MMM

[NCS90] Martin Neil, Robert J. Cole, and David Slater. Measures for maintenance management: A case study. Journal of Software Maintenance: Research and Practice, 2 (4):223–240, December 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Nandigam:1999:EEA

[NLČ99] Jagadeesh Nandigam, Arun Lakhotia, and Claude G. Čech. Experimental evaluation of agreement among programmers in applying the rules of cohesion. Journal of Software Maintenance: Research and Practice, 11(1):35– 53, January 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Nosek:1990:SMM

[NP90] John T. Nosek and Prashant Palvia. Software maintenance management: Changes in the last decade. Journal of Software Maintenance: Research and Practice, 2(3): 157–174, September 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Niessink:2000:SMS

[NvV00] Frank Niessink and Hans van Vliet. Software maintenance from a service perspective. Journal of Software Maintenance: Research and Practice, 12(2):103–120, March 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Newton:1992:RRC

[NYB92] Jenny Newton, Hongji Yang, and Keith Bennett. Report: Report on the Conference on Software Maintenance 1991, Sorrento, Italy, 14–17 October 1991. Journal of Software Maintenance: Research and Practice, 4(3): 163–175, September 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Olsem:1998:IAS

[Ols98] Michael R. Olsem. An incremental approach to software systems re-engineering. Journal of Software Maintenance: Research and Practice, 10(3):181–202, May 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

ONeill:1997:SMG

[O’N97] Don O’Neill. Software maintenance and global competitiveness. Journal of Software Maintenance: Research and Practice, 9(6):379–399, November 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Onoma:1995:SMI

[OTT^{+}95] A. K. Onoma, W. T. Tsai, F. Tsunoda, H. Suganuma, and S. Subramanian. Software maintenance — an industrial experience. Journal of Software Maintenance: Research and Practice, 7(5): 333–375, September 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Ogando:1994:OFP

[OYLW94] Roger M. Ogando, Stephen S.

Yau, Sying S. Liu, and Norman Wilde. An object finder for program structure understanding in software maintenance. Journal of Software Maintenance: Research and Practice, 6(5): 261–283, September 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Payne:1993:BRS

[Pay93]  Alun Payne.  Book re-

view: Software design techniques for large Ada systems, W. E. Byme. Digital Press (Prentice Hall, Hemel Hempstead). ISBN 0-13-816174-7. Price £39.00 (hardback). Journal of Software Maintenance: Research and Practice, 5(1):61–62, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Payne:1994:BRM

[Pay94a] Allen Payne. Book review:

Methods and tools for software configuration management, D. Whitgift. John Wiley & Sons, Chichester, 1991, ISBN 0-471-92940-9. Price: £29.95. Journal of Software Maintenance: Research and Practice, 6(2):102– 103, March 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Payne:1994:BRO

[Pay94b]  Allen Payne.  Book re-

view: Object oriented design, P. Robinson. Unicom Applied Technology, Chapman and Hall, Andover, 1992, ISBN 0-412-40520-2. Price: £65. Journal of Software Maintenance: Research and Practice, 6(2):103–104, March 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Papapanagiotakis:1994:SMM

[PB94] Giorgos Papapanagiotakis and Peter Breuer. A software maintenance management model based on queueing networks. Journal of Software Maintenance: Research and Practice, 6(2):73– 97, March 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Peercy:1997:IDP

[PC97] David E. Peercy and Ned Chapin. Interview with David E. Peercy. Journal of Software Maintenance: Research and Practice, 9(3):177– 200, May 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Pau:1990:SSM

[PK90] L. F. Pau and J. B. Kristinsson. SOFTM: A software maintenance expert system in Prolog. Journal of Software Maintenance: Research and Practice, 2(2):87– 111, June 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Pei:1994:RSM

[PV94] Gary Pei and Agustina Victoria. Reusability in software maintenance: A Japan– USA comparison. Journal of Software Maintenance: Research and Practice, 6(4):165– 183, July 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Rajlich:1997:MMS

[Raj97] VÁClav Rajlich. MSE: A methodology for software evolution. Journal of Software Maintenance: Research and Practice, 9(2):103–124, March 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Ricketts:1989:HDM

[Ric89] J. A. Ricketts. Harmonizing the development, maintenance and enhancement of application systems. Journal of Software Maintenance: Research and Practice, 1 (2):101–120, December 1989. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Rine:1991:ADI

[Rin91] David C. Rine. An approach for developing an integrated modelling environment: A case investigation. Journal of Software Maintenance: Research and Practice, 3(2):65–83, June 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Rine:1992:FAP

[Rin92] David Rine. A formal approach to perfective software maintenance: Integration using a basis. Journal of Software Maintenance: Research and Practice, 4(3): 119–128, September 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Rotenstreich:1990:ETD

[Rot90] Shmuel Rotenstreich. Enhancements through design transformations: A retroactive case study. Journal of Software Maintenance: Research and Practice, 2(4): 193–208, December 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Sakthivel:1994:DMC

[Sak94] Sachidanandam Sakthivel. A decision model to choose between software maintenance and software redevelopment. Journal of Software Maintenance: Research and Practice, 6(3):121–143, May 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Swanson:1989:OFM

[SB89] E. B. Swanson and C. M.

Beath. Organizational foundations for maintenance. Journal of Software Maintenance: Research and Practice, 1(1):47–58, September 1989. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Svahnberg:1999:ESP

[SB99] Mikael Svahnberg and Jan Bosch. Evolution in software product lines: two cases. Journal of Software Maintenance: Research and Practice, 11(6):391–422, November 1999. CODEN JSM-PEU. ISSN 1040-550X (print), 1096-908X (electronic).

Swanson:1995:IBS

[SC95] E. Burton Swanson and Ned Chapin. Interview with E. Burton Swanson. Journal of Software Maintenance: Research and Practice, 7(5): 303–315, September 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Scandura:1990:CAS

[Sca90] Joseph M. Scandura. Cognitive approach to systems engineering and re-engineering: Integrating new designs with old systems. Journal of Software Maintenance: Research and Practice, 2(3): 145–156, September 1990. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Schatzberg:1993:TQM

[Sch93] Doreen R. Schatzberg. Total quality management for maintenance process improvement. Journal of Software Maintenance: Research and Practice, 5(1):1–12, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Schach:1994:EIS

[Sch94] Stephen R. Schach. The economic impact of software reuse on maintenance. Journal of Software Maintenance: Research and Practice, 6(4):185–196, July 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Seaman:2000:BRP

[Sea00] Carolyn B. Seaman. Book review: Practical Software Maintenance. Journal of Software Maintenance: Research and Practice, 12(4): 249–253, July 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Segal:1989:DPU

[SF89] M. E. Segal and O. Frieder.

Dynamic program updating: A software maintenance technique for minimizing software, downtime. Journal of Software Maintenance: Research and Practice, 1(1):59–79, September 1989. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Sherer:1997:URA

[She97]  Susan A. Sherer.  Using

risk analysis to manage software maintenance. Journal of Software Maintenance: Research and Practice, 9 (6):345–364, November 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Sharpe:1991:CES

[SHH91] Shane Sharpe, Dwight A.

Haworth, and David Hale. Characteristics of empirical software maintenance studies: 1980–1989. Journal of Software Maintenance: Research and Practice, 3(1):1– 15, March 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Shepperd:1991:DMS

[SI91] Martin Shepperd and Darrel Ince. Design metrics and software maintainability: An experimental investigation.

Journal of Software Maintenance: Research and Practice, 3(4):215–232, December 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Sillince:1994:DIS

[Sil94] J. A. A. Sillince. A design for information systems which can adapt to changing organizational requirements. Journal of Software Maintenance: Research and Practice, 6(3):145–160, May 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Sneed:1994:DLA

[SN94] Harry M. Sneed and Erika Nyary. Downsizing large application programs. Journal of Software Maintenance: Research and Practice, 6(5): 235–247, September 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Sneed:1991:ESR

[Sne91] Harry M. Sneed. Economics of software re-engineering. Journal of Software Maintenance: Research and Practice, 3(3): 163–182, September 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Sneed:1995:UST

[Sne95] Harry M. Sneed. Understanding software through numbers: A metric based approach to program comprehension. Journal of Software Maintenance: Research and Practice, 7(6):405–419, November 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Stark:1995:SIU

[SO95] George E. Stark and Paul Oman. A survey instrument for understanding the complexity of software maintenance. Journal of Software Maintenance: Research and Practice, 7(6):421–441, November 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Stark:1997:SMM

[SO97] George E. Stark and Paul W.

Oman. Software maintenance management strategies: observations from the field. Journal of Software Maintenance: Research and Practice, 9(6):365–378, November 1997. CODEN JSM-PEU. ISSN 1040-550X (print), 1096-908X (electronic).

Stark:1999:EER

[SOSA99] George E. Stark, Paul Oman, Alan Skillicorn, and Alan Ameele. An examination of the effects of requirements changes on software maintenance releases. Journal of Software Maintenance: Research and Practice, 11 (5):293–309, September 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Schach:2000:MOA

[ST00] Stephen R. Schach and Amir Tomer. A maintenanceoriented approach to software construction. Journal of Software Maintenance: Research and Practice, 12(1):25– 45, January 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Sano:1997:ITS

[STR97] Takashi Sano, Wei-Tek Tsai, and Sanjai Rayadurgam. Interview with Takashi Sano. Journal of Software Maintenance: Research and Practice, 9(4):253–268, July 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Sere:1996:RED

[SW96] K. Sere and M. Waldén. Reverse engineering distributed algorithms. Journal of Software Maintenance: Research and Practice, 8(2):117–144, March 1996. CODEN JSM-PEU. ISSN 1040-550X (print), 1096-908X (electronic).

Tonella:1999:VPR

[TAFM99] P. Tonella, G. Antoniol, R. Fiutem, and E. Merlo. Variable-precision reaching definitions analysis. Journal of Software Maintenance: Research and Practice, 11 (2):117–142, March 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Tsalidis:1992:ASM

[TCM92] C. Tsalidis, D. Christodoulakis, and D. Maritsas. Athena: A software measurement and metrics environment. Journal of Software Maintenance: Research and Practice, 4(2): 61–81, June 1992. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Tan:1998:AMP

[TG98] Wui-Gee Tan and Guy G.

Gable. Attitudes of maintenance personnel towards maintenance work: a comparative analysis. Journal of Software Maintenance: Research and Practice, 10(1):59– 74, January 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Turver:1994:EIA

[TM94] Richard J. Turver and Malcolm Munro. An early impact analysis technique for software maintenance. Journal of Software Maintenance: Research and Practice, 6(1):35– 52, January 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Taylor:1998:TSM

[TML98] M. J. Taylor, E. P. Moynihan, and A. Laws. Training for software maintenance. Journal of Software Maintenance: Research and Practice, 10 (6):381–393, November 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Trapnell:1995:LEA

[Tra95] Fritz Trapnell. Letter to the Editor: Again, that maintenance ‘iceberg’. Journal of Software Maintenance: Research and Practice, 7(5):381, September 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Tsivkin:1996:LSP

[Tsi96] Vladimir M. Tsivkin. Live short and prosper: Applying cost analysis to routine software maintenance procedures. Journal of Software Maintenance: Research and Practice, 8(4):257–267, July 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Taramaa:1996:SCA

[TSM96] Jorma Taramaa, Veikko Seppänen, and Minna Mäkäräinen. From software configuration to application management

— improving the maturity of the maintenance of embedded software. Journal of Software Maintenance: Research and Practice, 8(1):49– 75, January 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Triantafyllos:1996:SMM

[TVDF96] George Triantafyllos, Stamatis Vassiliadis, and JosÉ G. Delgado-Frias. Software metrics and microcode: A case study. Journal of Software Maintenance: Research and Practice, 8(3):199–224, May 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Taylor:1996:MSM

[TWH96] M. J. Taylor and A. T. Wood-Harper. Methodologies and software maintenance. Journal of Software Maintenance: Research and Practice, 8(5): 295–307, September 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Vesely:1997:UCD

[Ves97] Eric Garrigue Vesely. Using COBOL defensive traps. Journal of Software Maintenance: Research and Practice, 9(5):329–342, September 1997. CODEN JSM-PEU. ISSN 1040-550X (print), 1096-908X (electronic).

VanDeursen:1998:LLL

[VK98] Arie Van Deursen and Paul Klint. Little languages: little maintenance? Journal of Software Maintenance: Research and Practice, 10(2):75– 92, March 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

vonMayrhauser:1997:PUB

[vMVH97] Anneliese von Mayrhauser, A. Marie Vans, and Adele E. Howe. Program understanding behaviour during enhancement of large-scale software. Journal of Software Maintenance: Research and Practice, 9(5):299–327, September 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

vonMayrhauser:1999:ART

[vMZ99] Anneliese von Mayrhauser and Ning Zhang. Automated regression testing using DBT and sleuth. Journal of Software Maintenance: Research and Practice, 11(2): 93–116, March 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Voas:1999:DIS

[Voa99] Jeffrey M. Voas. Disposable information systems: the future of software maintenance? Journal of Software Maintenance: Research and Practice, 11(2):143–150, March 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Vierimaa:1998:FTE

[VTP^{+}98] Matias Vierimaa, Jorma Taramaa, Heli Puustinen, Kati Suominen, and Tommi Ketola. Framework for tool evaluation for a maintenance environment. Journal of Software Maintenance: Research and Practice, 10(3): 203–224, May 1998. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

vanVliet:2000:RFN

[vV00] Hans van Vliet. Review of Frank Niessink’s thesis. Journal of Software Maintenance: Research and Practice, 12(3):185, May 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Ward:1993:ASC

[War93]  Martin Ward.  Abstract-

ing a specification from code. Journal of Software Maintenance: Research and Practice, 5(2):101–122, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Ward:1995:LEA

[War95a]  Dr. Martin Ward.  Let-

ter to the Editor: Abstraction refinement: A model of software evolution discussion by M. Ward. Journal of Software Maintenance: Research and Practice, 7(5): 377–378, September 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic). See [KN93].

Ward:1995:SCD

[War95b] Martin Ward. Short communication: A definition of abstraction. Journal of Software Maintenance: Research and Practice, 7(6): 443–450, November 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Ward:1995:FML

[WB95] M. P. Ward and K. H. Bennett. Formal methods for legacy systems. Journal of Software Maintenance: Research and Practice, 7(3):203– 219, May 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Wilde:1996:GPS

[WB96] Norman Wilde and Scott M. Brown. The GUMP process for software maintenance and maintenance education.

Journal of Software Maintenance: Research and Practice, 8(4):229–239, July 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

West:1996:BRI

[Wes96] Richard West. Book review:

Improving the Maintainability of Software. Journal of Software Maintenance: Research and Practice, 8(5): 345–356, September 1996. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Wendt:1997:SLM

[WI97] Cris Wendt and Nicholas Imparato. Software licensing models amid market turbulence. Journal of Software Maintenance: Research and Practice, 9(5):271–280, September 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Wilde:1999:DAM

[WJBW99] Norman Wilde, Randy Justice, Kristin Blackwell, and W. Eric Wong. Dynamic analysis methods for the year 2000 problem. Journal of Software Maintenance: Research and Practice, 11(3): 167–182, May 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Wild:1991:DBS

[WML91] Chris Wild, Kurt Maly, and Lianfang Liu. Decisionbased software development. Journal of Software Maintenance: Research and Practice, 3(1):17–43, March 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Welker:1997:PAD

[WOA97] Kurt D. Welker, Paul W. Oman, and Gerald G. Atkinson. Practice articles: Development and application of an automated source code maintainability index. Journal of Software Maintenance: Research and Practice, 9(3):127– 159, May 1997. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Wilde:1995:SRM

[WS95] Norman Wilde and Michael C. Scully. Software reconnaissance: Mapping program features to code. Journal of Software Maintenance: Research and Practice, 7(1):49– 62, January 1995. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Yate:1993:BRS

[Yat93] John Yate. Book review: Software sizing and estimating MKII FPA, Charles R. Symons. John Wiley & Sons (Chichester). ISBN 0-471-92985-9. Price £36.50 (hardback). Journal of Software Maintenance: Research and Practice, 5(1):59–61, ???? 1993. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Yang:2000:AKN

[YLZ00] Hongji Yang, Xiaodong Liu, and Hussein Zedan. Abstraction: a key notion for reverse engineering in a system reengineering approach. Journal of Software Maintenance: Research and Practice, 12(4):197–228, July 2000. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Yip:1991:WUI

[YR91] Stephen W. L. Yip and David J. Robson. Window user interfaces and software maintenance. Journal of Software Maintenance: Research and Practice, 3(2):107– 123, June 1991. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Younger:1994:IES

[YW94] E. J. Younger and M. P.

Ward. Inverse-engineering a simple real-time program. Journal of Software Maintenance: Research and Practice, 6(4):197–234, July 1994. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Zahran:1999:BRS

[Zah99]  Sami Zahran.  Book re-

view: Software process improvement: practical guidelines for business success, Addison Wesley Longman, UK, 1997, £27.95, 447 pp. Journal of Software Maintenance: Research and Practice, 11(4):285–291, July 1999. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).

Zimmer:1989:AC

[Zim89]  J. A. Zimmer.  Altering

without change. Journal of Software Maintenance: Research and Practice, 1(1):29– 46, September 1989. CODEN JSMPEU. ISSN 1040-550X (print), 1096-908X (electronic).